### v2

@@ 3.4 Transfer operator for a nonlinearly perturbed Arnold’s cat map
We consider a nonlinear perturbation of the (linear and hyperbolic) Arnold’s cat map of the two-torus $X=\mathbb{T}^{2}$ introduced in [9]:
@@
with $\delta=0.01$. With this perturbation, $T$ remains Anosov and therefore has an SRB measure $\mu$ [5, Theorem 4.12]. The spectral theory for Anosov diffeomorphisms is challenging and relies on carefully constructed anisotropic Banach spaces (e.g. [15]) that are adapted to the unstable and stable foliations. In this example, we let $\mathcal{L}f=f\circ T^{-1}/|\det DT\circ T^{-1}|$ be the Perron–Frobenius operator, and consider $\mathcal{L}:L^{2}(X,m)\to L^{2}(X,m)$ where $m$ is Lebesgue measure on the flat torus. Note the additional determinant term in the definition of $\mathcal{L}$; the constant function is no longer the leading eigenfunction of $\mathcal{L}$. In fact, the object $\mu$ satisfying $\mathcal{L}\mu=\mu$ is in general a distribution and need not lie in $L^{2}(\mathbb{T}^{2},m)$. The fixed distribution $\mu$ may be identified with the SRB measure and describes the long-term distribution of infinitely long trajectories of $T$ for Lebesgue almost all initial conditions $x\in\mathbb{T}^{2}$. This distribution of infinitely long trajectories is accessible through the fixed point of $\mathcal{L}$; this is one of the many advantages of studying transfer operators.
[…]
Figure 8: Input and output for the Perron–Frobenius operator of a perturbed Arnold’s cat map. Left to right, the three panels show (i) an unseen input observable $g$, (ii) the corresponding ground-truth output from the true transfer operator, (iii) and the corresponding predicted output from our learned transfer operator. The mean $L^{2}$ relative error over the full test set is $1.225\times 10^{-2}$.
Next, we analyse the learned set of basis functions $\{\phi_{j}\}_{j=1}^{N}$ that span the invariant subspace $V$. Fig. 9 shows representative basis functions obtained without sparsity penalty. The basis functions obtained through the proposed network show unmistakable alignment with contracting and expanding directions of the nonlinear dynamics generated by $T$.
Figure 9: Representative basis functions learned by the network. The patterns exhibit structures that align with the map’s contracting and expanding directions, illustrating how the learned invariant subspace is adapted to the underlying dynamics.
@@
Figure 11: Eigenvalues of the learned Perron–Frobenius operator for the model with no sparsity penalty. The leading eigenvalue (corresponding to the SRB measure) is close to 1 as expected. The remaining eigenvalues are separated away from the unit circle.
The remaining eigenvalues in Fig. 11 are restricted to a disk away from the unit circle, in line with expectations if one used a suitable anisotropic Banach space for the domain of $\mathcal{L}$ [15]. While we work in $L^{2}(X,m)$, we train our estimate of $\mathcal{L}$ on smooth functions, and it is likely that this restriction to smooth functions produces the spectral gap we see in Fig. 11.
@@
Figure 12: Densities of ground-truth SRB measure (left), its reconstruction using the SABON learned basis (middle), and its reconstruction using a Fourier basis of the same dimension (right). We consider the densities as vectors on the grid and make their $\ell_{1}-$norms equal to 1.
@@
To quantitatively benchmark the learned basis against the traditional Fourier basis we measure both the relative $L^{2}$ error and Wasserstein $W_{1}$ distance of the orthogonally projected ground-truth SRB measure onto each basis. Letting $\mathcal{B}$ denote either of the two bases, denote the orthogonal projector onto $\mathcal{B}$ by $P_{\mathcal{B}}:L^{2}(X)\to\operatorname{span}(\mathcal{B})$. Define the relative $L^{2}$ projection error
$L^{2,\mathrm{rel}}_{\mathcal{B}}\bigl{(}\rho_{\mathrm{SRB}},P_{\mathcal{B}}(\rho_{\mathrm{SRB}})\bigr{)}\;\coloneq\;\frac{\|\rho_{\mathrm{SRB}}-P_{\mathcal{B}}(\rho_{\mathrm{SRB}})\|_{2}}{\|\rho_{\mathrm{SRB}}\|_{2}}.$
An SRB measure is typically a distribution and not a function, and therefore it has no density with respect to Lebesgue measure. This makes comparisons in the $L^{2}$ norm less meaningful (the $L^{2}$ topology is too strong). More appropriate is the 111-Wasserstein metric on probability measures. The 1-Wasserstein distance is defined as
$W_{1}(\rho,\eta)\;\coloneq\;\inf_{\pi\in\Pi(\rho,\eta)}\int_{X\times X}d(x,y)\,\mathrm{d}\pi(x,y),$
where $d(x,y)$ is the metric distance between $x$ and $y$ on the phase space $X$ and $\Pi(\rho,\eta)$ denotes all probability measures $\pi$ on $X\times X$ whose first marginal is $\rho$ and whose second marginal is $\eta$. After discretising $\rho$ and $\eta$ onto the uniform grid $\{x_{i}\}_{i=1,\dots,n}$, the infimum above can be written as the linear program
$\min_{\pi\geq 0}\;\sum_{i,j=1}^{n}\pi_{ij}C_{ij}\quad\text{s.t.}\quad\sum_{j=1}^{n}\pi_{ij}=\rho_{i},\;\sum_{i=1}^{n}\pi_{ij}=\eta_{j},$
where $C_{ij}$ is the distance between $x_{i}$ and $x_{j}$ in $X$ for $i,j=1,\ldots,n$, $\pi_{ij}$ is the mass flow between $x_{i}$ and $x_{j}$, and $0\leq\rho_{i},\eta_{i}\leq 1$ are the discrete masses assigned by $\rho,\eta$ (respectively) to the point $x_{i}\in X$ for each $i=1,\ldots,n$.
@@
As in the relative $L^{2}$ norm calculation, we compute $W_{1}\bigl{(}\rho_{\mathrm{SRB}},P_{\mathcal{B}}(\rho_{\mathrm{SRB}})\bigr{)}$ when (i) $\mathcal{B}$ is the learned basis and (ii) $\mathcal{B}$ is a Fourier basis of comparable cardinality. Table 5 shows that in both $L^{2}$ and Wasserstein distances, the learned SABON basis achieves the smallest error, which implies that the learned basis is able to better capture the detailed structure in the “ground truth” SRB distribution.
Table 5: Approximation error of the ground-truth SRB measure under orthogonal projection onto the two candidate bases. Lower values denote a more accurate representation.
Basis $L^{2,\mathrm{rel}}_{\mathcal{B}}\bigl{(}\rho_{\mathrm{SRB}},P_{\mathcal{B}}(\rho_{\mathrm{SRB}})\bigr{)}$ $W_{1}\bigl{(}\rho_{\mathrm{SRB}},P_{\mathcal{B}}(\rho_{\mathrm{SRB}})\bigr{)}$
SABON (225 functions) $7.479\times 10^{-2}$ $7.886\times 10^{-4}$
Fourier (225 functions) $1.318\times 10^{-1}$ $1.717\times 10^{-3}$
Finally, we quantify and compare the invariance of the subspaces generated by the learned basis and a classical Fourier basis of the same size under the action of the Perron–Frobenius operator. We compute principal angles between each subspace and its image under the operator, summarising these in Table 6. The significantly smaller distances obtained for our learned subspace indicate superior invariance compared with the traditional Fourier basis.
Table 6: Subspace‐distance measures quantifying how well the subspace is preserved under the Perron–Frobenius operator, comparing the SABON basis (225 elements, no sparsity penalty) with a Fourier basis with 225 functions. Lower values indicate greater invariance.
Metric SABON basis Fourier basis
Geodesic ($d_{\max}$) 0.0370.0370.037 16.64216.64216.642
Chordal ($d_{\text{ch}}$) 0.0370.0370.037 10.61510.61510.615
Projection ($d_{\text{pr}}$) 0.0260.0260.026 1.0001.0001.000
A more detailed comparison is shown graphically in Fig. 13, where we plot the projection distance for every individual principal angle for both subspaces.
Figure 13: Projection distance for the principal angles between each subspace and its image under the Perron–Frobenius operator for the learned basis and the Fourier basis of equal cardinality. Smaller distances correspond to greater invariance.
@@
Why does the SABON basis have far superior invariance when compared to a Fourier basis? A Fourier element — whether the complex exponential $e^{2\pi ik\cdot x}$ or the purely real sine–cosine products introduced in Section 3.1 — is aligned with coordinate axes. The contracting dynamics of $T$ means that the frequency of a Fourier mode will be increased in stable directions of the dynamics under application of $\mathcal{L}$. Thus, those modes in the Fourier basis with higher frequencies may be pushed far out of the original span of the Fourier modes, and made near-orthogonal to the Fourier basis. Indeed, Fig. 13 indicates that this occurs to more than half of the Fourier modes.
@@
The SABON basis, on the other hand, is learned jointly with the dynamics, so the network discovers functions that organise along the stable and unstable directions. The resulting modes resemble elongated ridges: smooth in the expanding direction and more sharply confined in the contracting one. When such a ridge is pushed forward by $\mathcal{L}$ it tends to become another ridge, consistent with the invariance of the SABON subspace. This anisotropic feature, absent from the isotropic Fourier family, accounts for the much smaller distances reported for SABON in Table 6 and illustrated in Fig. 13.
@@
A further crucial aspect of our learned SABON basis is its efficiency. As mentioned above, our learned basis functions tend to align with the stable and unstable directions arising from the dynamics. This yields basis functions that have more rapid oscillation in stable directions and slower oscillation in unstable directions, which is exactly the form one expects for eigendistributions [15]. With a comparable number of basis functions, our learned functions can therefore more accurately represent the eigenprojectors.

### v3

@@ 3.4.1 A nonlinearly perturbed Arnold’s cat map
We consider a nonlinear perturbation of the (linear and hyperbolic) Arnold’s cat map of the two-torus $X=\mathbb{T}^{2}$ introduced in [CF20]:
@@
with $\delta=0.01$. Under this perturbation, $T$ remains Anosov and therefore has an SRB measure $\mu_{\mathrm{SRB}}$ [bowen1975, Theorem 4.12]. The spectral theory for Anosov diffeomorphisms is challenging and relies on carefully constructed anisotropic Banach spaces (e.g. [GL06]) that are adapted to the unstable and stable foliations. In this example, we let $\mathcal{L}f=f\circ T^{-1}/|\det DT\circ T^{-1}|$ be the Perron–Frobenius operator, and consider $\mathcal{L}:L^{2}(X,m)\to L^{2}(X,m)$ where $m$ is Lebesgue measure on the flat torus. Note the additional determinant term in the definition of $\mathcal{L}$; the constant function is no longer the leading eigenfunction of $\mathcal{L}$. In fact, the object $\mu_{\mathrm{SRB}}$ satisfying $\mathcal{L}\mu_{\mathrm{SRB}}=\mu_{\mathrm{SRB}}$ is in general a distribution and need not lie in $L^{2}(\mathbb{T}^{2},m)$. The fixed distribution $\mu_{\mathrm{SRB}}$ may be identified with the SRB measure and describes the long-term distribution of infinitely long trajectories of $T$ for Lebesgue almost all initial conditions $x\in\mathbb{T}^{2}$. This distribution of infinitely long trajectories is accessible through the fixed point of $\mathcal{L}$; this is one of the many advantages of studying transfer operators.
To eliminate the discontinuities that arise at the edges of the unit square, we embed $\mathbb{T}^{2}$ into $\mathbb{R}^{4}$ by mapping each $(x,y)$ to its pair of complex phases $(\cos 2\pi x,\allowbreak\sin 2\pi x,\allowbreak\cos 2\pi y,\allowbreak\sin 2\pi y)$. That is, we treat a problem with intrinsic dimension $d^{\prime}=2$ in a higher ambient space with dimension $d=4$. We first evaluate our model’s ability to approximate the Perron–Frobenius operator on an unseen (not in the training set) observable. To highlight the anisotropic nature of the learned basis everywhere in the phase space, we chose not to apply the sparsity penalty throughout both Anosov examples. Fig. 8 displays an unseen input observable $g$, the ground-truth image, and our network’s prediction. The accuracy is quantified by the relative $L^{2}$ error of $1.625\times 10^{-2}$ on the full test set.
Figure 8: Input and output for the Perron–Frobenius operator of a perturbed Arnold’s cat map. Left to right, the three panels show (i) an unseen input observable $g$, (ii) the corresponding ground-truth output from the true transfer operator, (iii) and the corresponding predicted output from $\hat{\mathcal{L}}$. The mean $L^{2}$ relative error over the full test set is $1.625\times 10^{-2}$.
Next, we analyse the learned set of basis functions $\{\phi_{j}\}_{j=1}^{N}$ that span the subspace $V$. Fig. 9 shows representative basis functions obtained without sparsity penalty. The basis functions obtained through the proposed network show unmistakable alignment with contracting and expanding directions of the nonlinear dynamics generated by $T$.
Figure 9: Representative basis functions learned by the network. The patterns exhibit structures that align with the map’s contracting and expanding directions, illustrating how the learned subspace is adapted to the underlying dynamics.
@@
Figure 11: Eigenvalues of $\hat{\mathcal{L}}$ for the model with no sparsity penalty. The leading eigenvalue (corresponding to the SRB measure) is close to 1 as expected. The remaining eigenvalues are separated away from the unit circle.
The remaining eigenvalues in Fig. 11 are restricted to a disk away from the unit circle, in line with expectations if one used a suitable anisotropic Banach space for the domain of $\mathcal{L}$ [GL06]. While we work in $L^{2}(X,m)$, we train our estimate of $\mathcal{L}$ on smooth functions, and it is likely that this restriction to smooth functions produces the spectral gap we see in Fig. 11.
@@
Figure 12: Densities of ground-truth SRB measure (left), its approximation using the SABON basis (middle), and using a Fourier basis of the same cardinality (right). In both cases, we use only the respective basis functions together with the true dynamics to solve the resulting eigenproblem, ensuring a fair comparison of basis quality. The densities are normalised to have unit $L^{1}(m)$ norm.
The “ground truth” SRB measure is constructed using a Galerkin method with $N=100$ complex Fourier modes per coordinate, yielding $N^{2}=10,000$ basis functions in total. The transfer operator matrix entries are computed via FFT on a fine $4N\times 4N=400\times 400$ spatial grid to ensure accurate quadrature. The leading eigenfunction is then evaluated on the $100\times 100$ grid used throughout our experiments. In comparison, our approximated SRB measure uses 324 learned basis functions. We also compare against a Fourier basis of the same cardinality. We use the real Fourier basis on $0,1]^{2}$ constructed as a tensor product of 1D bases with 18 functions each: one constant, sine-cosine pairs for modes of order $1,\ldots,8$, and a single cosine at order $9$ (the corresponding sine is omitted to obtain exactly 324 functions for a fair comparison). The tensor product yields $18^{2}=324$ basis functions in total. From Figure 12 we can see that our reconstructed SRB measure using our learned basis clearly captures more detail and relevant dynamical structure than a high-quality basis created from a similar number of Fourier modes.
An SRB measure is in general a distribution and not a function, and therefore it has no density with respect to Lebesgue measure. This makes comparisons in the $L^{2}$ norm less meaningful. Negative index Sobolev spaces [MMP05, Thiffeault12] have been used to describe chaotic mixing because they contain distributions of certain orders and their norms are simple to evaluate. We choose the space $H^{-1}$ as advocated in [Thiffeault12] because it is in spirit compatible with the lowest-order Banach space from [GL06]. For a function $f$ on the torus $\mathbb{T}^{d}$ with Fourier coefficients $\hat{f}_{k}$, the $H^{-1}$ norm is defined as
$\|f\|_{H^{-1}}^{2}\;\coloneq\;\sum_{k\in\mathbb{Z}^{d}}(1+|k|^{2})^{-1}|\hat{f}_{k}|^{2},$
where $|k|=\sqrt{k_{1}^{2}+\cdots+k_{d}^{2}}$ is the Euclidean norm of the multi-index $k$. Using the above expression, the $H^{-1}$ norm is straightforward to evaluate via the FFT. This norm penalises high-frequency components less severely than $L^{2}$, making it natural for comparing distributions that may differ in fine-scale structure.
@@
To quantitatively benchmark the learned basis against the traditional Fourier basis, we measure two types of error. First, the projection error, which will quantify how well each basis can represent the ground-truth SRB measure under both $L^{2}$ and $H^{-1}$ orthogonal projection in their respective norms. Second, the approximation error compares the estimated SRB measure $\mu_{\mathcal{B}}$ constructed as an eigenvector of $\mathcal{L}_{\mathcal{B}}$ (Galerkin projection of $\mathcal{L}$ onto the learned basis $\mathcal{B}$) against the ground truth in both the $L^{2}$ and $H^{-1}$ norms.
@@
Letting $\mathcal{B}$ denote either of the two bases, denote the $L^{2}$-orthogonal projector onto $\mathcal{B}$ by $P_{\mathcal{B}}:L^{2}(X)\to\operatorname{span}(\mathcal{B})$. Define the relative $L^{2}$ projection error
(10) $L^{2,\mathrm{rel}}_{\mathcal{B}}\bigl(\mu_{\mathrm{SRB}},P_{\mathcal{B}}(\mu_{\mathrm{SRB}})\bigr)\;\coloneq\;\frac{\|\mu_{\mathrm{SRB}}-P_{\mathcal{B}}(\mu_{\mathrm{SRB}})\|_{2}}{\|\mu_{\mathrm{SRB}}\|_{2}},$
and define the relative $L^{2}$ approximation error for $\mu_{B}$
(11) $L^{2,\mathrm{rel}}\bigl(\mu_{\mathrm{SRB}},{\mu}_{\mathcal{B}}\bigr)\;\coloneq\;\frac{\|\mu_{\mathrm{SRB}}-{\mu}_{\mathcal{B}}\|_{2}}{\|\mu_{\mathrm{SRB}}\|_{2}},$
We define the relative $H^{-1}$ projection error in an identical way to the $L^{2}$ projection error, except the projection is $H^{-1}$-orthogonal and the norm is $H^{-1}$:
(12) $H^{-1,\mathrm{rel}}_{\mathcal{B}}\bigl(\mu_{\mathrm{SRB}},P_{\mathcal{B}}(\mu_{\mathrm{SRB}})\bigr)\;\coloneq\;\frac{\|\mu_{\mathrm{SRB}}-P_{\mathcal{B}}(\mu_{\mathrm{SRB}})\|_{H^{-1}}}{\|\mu_{\mathrm{SRB}}\|_{H^{-1}}}.$
The relative $H^{-1}$ approximation error for $\mu_{\mathcal{B}}$ is
(13) $H^{-1,\mathrm{rel}}\bigl(\mu_{\mathrm{SRB}},{\mu}_{\mathcal{B}}\bigr)\;\coloneq\;\frac{\|\mu_{\mathrm{SRB}}-{\mu}_{\mathcal{B}}\|_{H^{-1}}}{\|\mu_{\mathrm{SRB}}\|_{H^{-1}}}.$
@@
We compute these quantities when (i) $\mathcal{B}$ is the learned basis and (ii) $\mathcal{B}$ is a Fourier basis of comparable cardinality. Table 4 shows that in both relative $L^{2}$ and $H^{-1}$ norms, the learned SABON basis has a consistently smaller error (by a factor of around four) for both projection and approximation, further illustrating that the learned basis is able to better capture the detailed structure in the ground truth SRB distribution.
Table 4: Approximation quality of the ground-truth SRB measure for the weakly nonlinear cat map. Projection error (Eqs. 10 and 12) measures how well each basis can represent the SRB measure under orthogonal projection. Approximation error (Eqs. 11 and 13) compares the estimated SRB measure $\mu_{\mathcal{B}}$, computed by solving ${\mathcal{L}_{\mathcal{B}}}\mu_{\mathcal{B}}=\mu_{\mathcal{B}}$, with the ground-truth SRB measure.
Projection Error Approximation Error
Basis ($N=324$) $L^{2,\mathrm{rel}}$ $H^{-1,\mathrm{rel}}$ $L^{2,\mathrm{rel}}$ $H^{-1,\mathrm{rel}}$
SABON $6.436\times 10^{-2}$ $1.016\times 10^{-3}$ $7.314\times 10^{-2}$ $4.064\times 10^{-3}$
Fourier $2.647\times 10^{-1}$ $1.576\times 10^{-2}$ $2.648\times 10^{-1}$ $1.578\times 10^{-2}$
3.4.2 A strongly conjugated nonlinear perturbation of Arnold’s cat map
To test SABON on dynamics with more complex geometry, we consider a smooth conjugation of the perturbed cat map from Section 3.4.1. Let $F:\mathbb{T}^{2}\to\mathbb{T}^{2}$ be the diffeomorphism
$F(x,y)=\bigl(x-a\sin(2\pi x),\;y+b\sin(2\pi y+\pi/4)\bigr)\pmod{1}$,
with $a=b=0.1$. We define the conjugated map
$T_{\mathrm{conj}}\coloneq F^{-1}\circ T\circ F,$
where $T$ is the perturbed cat map from Section 3.4.1 with $\delta=0.01$. Since $T_{\mathrm{conj}}$ is diffeomorphically conjugate to $T$, it remains Anosov and possesses an SRB measure $\mu_{\rm SRB_{\mathrm{conj}}}=(F^{-1})_{*}\mu_{\rm SRB}=\mu_{\rm SRB}\circ F$, where $\mu_{\rm SRB}$ is the SRB measure of $T$.
The transfer operator for $T_{\mathrm{conj}}$ is given by
$\mathcal{L}_{\mathrm{conj}}f=\frac{f\circ T_{\mathrm{conj}}^{-1}}{|\det DT_{\mathrm{conj}}\circ T_{\mathrm{conj}}^{-1}|},$
where the Jacobian determinant of $T_{\mathrm{conj}}$ at a point $z$ satisfies (by the chain rule and the inverse function theorem)
$|\det DT_{\mathrm{conj}}(z)|=\frac{|\det DT(F(z))|\cdot|\det DF(z)|}{|\det DF(T_{\mathrm{conj}}(z))|}.$
We train SABON on $T_{\rm conj}$ using the same architecture and training procedure as in Section 3.4.1, but with a larger basis of $N=676$ functions to accommodate the more complex phase space geometry. For comparison, we use a Fourier basis constructed as in the previous example with 26 modes per dimension (maximum order 13), yielding $26^{2}=676$ functions. Fig. 13 displays the model’s performance on an unseen observable.
Figure 13: Input and output for the Perron–Frobenius operator of the conjugated cat map. Left to right: (i) an unseen input observable $g$, (ii) ground-truth output, (iii) predicted output. The mean $L^{2}$ relative error over the test set is $1.736\times 10^{-2}$.
The learned basis functions, shown in Fig. 14, exhibit clear adaptation to the geometry of the stable and unstable directions. As in the previous examples, the learned basis is approximately orthogonal; we omit the Gram matrix for brevity. Unlike the weakly nonlinear case in Section 3.4.1, where the basis functions display nearly linear striations approximately aligned with the linear cat map’s eigendirections, the basis functions here show pronounced curvature reflecting the distortion introduced by the conjugacy $F$. This demonstrates SABON’s ability to automatically discover basis functions adapted to the underlying dynamical geometry, even when that geometry is significantly more complex than the linear case.
Figure 14: Representative basis functions learned for the conjugated cat map. The curved structures reflect the nonlinear geometry of the stable and unstable directions induced by the conjugacy.
Solving the eigenequation (9) as in the previous examples, Fig. 15 shows the numerical eigenvalues of $\hat{\mathcal{L}}$. Despite the increased geometric complexity of the conjugated map, the spectral structure remains well-resolved: the leading eigenvalue lies close to unity, confirming the presence of an invariant measure, while the remaining eigenvalues are contained within a disk away from the unit circle.
To compare the quality of the SABON and Fourier bases independently of the learned operator, we compute the SRB measure as the leading eigenvector of the transfer operator matrix constructed using the true dynamics in both cases. That is, for each basis we project the true transfer operator onto the respective subspace and solve the resulting eigenequation. Fig. 16 shows that the SABON basis captures the fine-scale structure of the invariant measure more accurately than the Fourier basis, which provides a smoother approximation with less detail in the curved features induced by the conjugacy.
Figure 15: Eigenvalues of the learned Perron–Frobenius operator for the strongly nonlinear conjugated cat map. As in the weakly nonlinear case, the leading eigenvalue (corresponding to the SRB measure) is close to 1, and the remaining eigenvalues exhibit a spectral gap.
Figure 16: Densities of ground-truth SRB measure for the conjugated map (left), its approximation using the SABON basis (middle), and using a Fourier basis of the same cardinality (right). In both cases, we use only the respective basis functions together with the true dynamics to solve the resulting eigenproblem, ensuring a fair comparison of basis quality. The densities are normalised to have unit $L^{1}(m)$ norm.
Following the same procedure as in Section 3.4.1, we report the projection and approximation errors in Table 5.
Table 5: Approximation quality of the ground-truth SRB measure for the strongly nonlinear conjugated cat map. Projection error (Eqs. 10 and 12) measures how well each basis can represent the SRB measure under orthogonal projection. Approximation error (Eqs. 11 and 13) compares the estimated SRB measure $\mu_{\mathcal{B}}$, computed by solving ${\mathcal{L}_{\mathcal{B}}}\mu_{\mathcal{B}}=\mu_{\mathcal{B}}$, with the ground-truth SRB measure.
Projection Error Approximation Error
Basis ($N=676$) $L^{2,\mathrm{rel}}$ $H^{-1,\mathrm{rel}}$ $L^{2,\mathrm{rel}}$ $H^{-1,\mathrm{rel}}$
SABON $6.091\times 10^{-2}$ $466\times 10^{-3} 1.1$$147\times 10^{-1}$ $6.225\times 10^{-3}$
Fourier $2.409\times 10^{-1}$ $1.095\times 10^{-2}$ $2.414\times 10^{-1}$ $1.132\times 10^{-2}$
@@ 3.4.2 A strongly conjugated nonlinear perturbation of Arnold’s cat map
Fourier modes have many strong approximation properties and are well suited to the periodic domains we have considered. Why then does the SABON basis consistently produce more accurate estimates of $\mu_{\rm SBR}$ than a Fourier basis? Firstly, a Fourier element — whether the complex exponential $e^{2\pi ik\cdot x}$ or the purely real sine–cosine products introduced in Section 3.1 — is aligned with coordinate axes. Secondly, the contracting dynamics of $T$ means that the frequency of a Fourier mode will be increased in stable directions of the dynamics under application of $\mathcal{L}$. Thus, those modes in the Fourier basis with higher frequencies may be pushed far out of the original span of the Fourier modes, and made near-orthogonal to the Fourier basis.
@@
The SABON basis, on the other hand, is learned jointly with the dynamics, so the network discovers functions that organise along the stable and unstable directions. The resulting modes resemble elongated ridges: smooth in the expanding direction and more sharply confined in the contracting one. When such a ridge is pushed forward by $\mathcal{L}$ it tends to become another ridge. This anisotropic feature, absent from the isotropic Fourier family, accounts for the superior approximation quality reported in Table 4 and Table 5.
@@
A further crucial aspect of our learned SABON basis is its efficiency. As mentioned above, our learned basis functions tend to align with the stable and unstable directions arising from the dynamics. This yields basis functions that have more rapid oscillation in stable directions and slower oscillation in unstable directions, which is exactly the form one expects for eigendistributions [GL06]. With a comparable number of basis functions, our learned functions can therefore more accurately represent the eigenprojectors.